The length of a rectangle is 5 meters longer than the width. Find the length and width if the perimeter is 54 meters.
Answer by mr.barrett(5)

In a rectangle, the following equation relates the perimeter, length, and width:
Perimeter = 2*(length) + 2*(width).


In this rectangle,
Perimeter = 54
Width = w
Length = w + 5


By substituting, we have 54 = 2*(w+5) + 2*(w).
See if you can solve that equation for w.